oidjviq:

Part 1. A chemist reacted 12.0 liters of F2 gas with NaCl in the laboratory to form Cl2 gas and NaF. Use the ideal gas law equation to determine the mass of NaCl that reacted with F2 at 280. K and 1.50 atm. F2 + 2NaCl → Cl2 + 2NaF Part 2. Explain how you would determine the mass of sodium chloride that can react with the same volume of fluorine gas at STP. (10 points)

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie Is this a test question? \(\color{#0cbb34}{\text{End of Quote}}\) Also, This was posted in the wrong subject, next time make sure it's in the right subject.

toga:

Part 1: To determine the mass of NaCl that reacted with F2 gas, we can use the ideal gas law equation: PV = nRT where P is the pressure, V is the volume, n is the number of moles, R is the gas constant, and T is the temperature in Kelvin. First, we need to find the number of moles of F2 gas that reacted. We can do this using the equation: n = PV/RT where P = 1.50 atm, V = 12.0 L, R = 0.08206 L.atm/mol.K, and T = 280 K. n = (1.50 atm x 12.0 L) / (0.08206 L.atm/mol.K x 280 K) = 0.719 mol Since the balanced equation shows that 1 mole of F2 reacts with 2 moles of NaCl, we can determine the number of moles of NaCl that reacted as: n(NaCl) = 0.719 mol / 2 = 0.3595 mol Now, we can use the molar mass of NaCl to calculate the mass of NaCl that reacted: mass(NaCl) = n(NaCl) x M(NaCl) where M(NaCl) = 58.44 g/mol is the molar mass of NaCl. mass(NaCl) = 0.3595 mol x 58.44 g/mol = 21.01 g Therefore, the mass of NaCl that reacted with F2 gas is 21.01 g. Part 2: To determine the mass of sodium chloride that can react with the same volume of fluorine gas at STP, we need to use the ideal gas law equation again. At STP (Standard Temperature and Pressure), the temperature is 273 K and the pressure is 1 atm. Therefore, the ideal gas law equation can be written as: PV = nRT where P = 1 atm, V = 22.4 L (molar volume of any gas at STP), R = 0.08206 L.atm/mol.K, and T = 273 K. Now, we can use this equation to determine the number of moles of F2 gas present in 22.4 L at STP: n = PV/RT = (1 atm x 22.4 L)/(0.08206 L.atm/mol.K x 273 K) = 1.00 mol Since 1 mole of F2 gas reacts with 2 moles of NaCl, we can determine the number of moles of NaCl that can react as: n(NaCl) = 1.00 mol / 2 = 0.50 mol Finally, we can use the molar mass of NaCl to calculate the mass of NaCl that can react: mass(NaCl) = n(NaCl) x M(NaCl) = 0.50 mol x 58.44 g/mol = 29.22 g Therefore, the mass of NaCl that can react with the same volume of fluorine gas at STP is 29.22 g.